# Satellite Data Shows Resilience of Tigrayan Farmers in Crop Cultivation During Civil War

### v1

@@ Abstract
The Tigray War was an armed conflict that took place primarily in the Tigray region of northern Ethiopia from November 3, 2020 to November 2, 2022. Given the importance of agriculture in Tigray to livelihoods and food security, determining the impact of the war on cultivated area is critical, but quantifying this impact was difficult due to restricted movement within and into the region due to conflict-driven insecurity and blockages. Using satellite imagery and statistical area estimation techniques, we assessed changes in crop cultivation area in Tigray before and during the war. Our findings show that cultivated area was largely stable between 2020-2021 despite the widespread impacts of the war. We estimated $1,132,000\pm 133,000$ hectares of cultivation in pre-war 2020 compared to $1,217,000\pm 132,000$ hectares in mid-war 2021. Comparing changes inside and outside of a 5 km buffer around conflict events, we found a slightly higher upper confidence limit of cropland loss within the buffer (0-3%) compared to outside the buffer (0-1%). Our results support other reports that despite widespread war-related disruptions, Tigrayan farmers were largely able to sustain cultivation. Our study demonstrates the capability of remote sensing combined with machine learning and statistical techniques to provide timely, transparent area estimates for monitoring food security in regions inaccessible due to conflict.
@@ 1 Introduction
Agriculture is a critical industry in Tigray, providing employment for 80% of people and constituting 65% of the land area [6]. Disruptions to cultivation could have severe consequences for livelihoods and food security, thus quantifying the impact of the war on agriculture is crucial for targeting humanitarian relief to ensure the population’s basic needs are met. Many farmers reported delays in ploughing due to insecurity [3] and 353,000 people in Tigray were reportedly experiencing “catastrophic” levels of “acute food insecurity” in June 2021 [6]. Food insecurity in early 2021 was exacerbated by the loss of food stocks destroyed by armed forces during the start of the war in late 2020 and the blockage of food and humanitarian aid to the region [6]. Humanitarian aid was blocked or severely limited from the start of the conflict until April 2021, when aid organizations could deliver food aid for the first time since late 2020. In surveys conducted between April and May 2021, 50-70% of households in NorthWestern and Central Tigray reported that food assistance was their primary source of food [6]. The June 2021 cease-fire gave hope that farmers would be able to cultivate their fields in the 2021 main season, but there were grave concerns and uncertainty about food insecurity and the fate of food production in 2021.
Satellite remote sensing data can provide qualitative and quantitative insights about the impact of conflict on agricultural land use and production. This remote, real-time data source is especially critical for assessing impacts in regions where traditional ground-based surveys cannot be conducted due to insecurity, such as Tigray during 2021. Most studies using satellite data to assess war impacts on cultivation in other regions have found significant reductions in cultivated land due to war, ranging from approximately 15-30% across studies. Skakun et al. [7] found a 22% net loss of cropland in regions not occupied by the Ukraine government following the military conflict that began in 2014 in southeastern Ukraine, and that cropland losses were higher (46%) in areas in the buffer zone surrounding the conflict border. In a study of the more recent war in Ukraine that began in 2022, Ma et al. [8] detected concentrations of fallowed croplands in western Kherson and central Luhansk associated with conflict activities. Witmer [9] found that the average cultivated area decreased by 25% in the period after the Bosnian War between 1992-1995. In a study of the armed conflict in Mali, Boudinaud and Orenstein [10] found that 25% of areas in Mopti exhibited crop area losses in 2019 compared to pre-conflict years. Olsen et al. [11] concluded that cropland cultivation decreased by 16% from 2016 to 2018 as a result of armed conflict in South Sudan. Müller and Kummerle [12] and Peterson and Aunap [13] studied the impact of government and economic transitions in Romania after 1989 and Estonia between 1990-1993, respectively; while these transitions were not armed conflicts, they were tumultuous periods with disruptions to the agricultural sector that were associated with 32% of croplands being abandoned in Estonia and 21% abandonment in Romania. In the first (1994-1996) and second (1999-2009) Chechen wars, Yin et al. [14] found higher rates of cropland abandonment in areas closer to conflict events (45% abandonment within 1 km compared to 6% greater than 8 km from conflict events) as well as higher rates of re-cultivation in areas farther away from conflict events. While most studies have concluded net losses or negative impacts to agricultural production due to conflict, Li et al. [15] found that there was a net increase in rural areas of Syria during the civil war and Eklund et al. [16] reported a mixture of cropland expansion, abandonment, and other changes in areas of Syria and Iraq occupied by the Islamic State. NASA Harvest [17] found that despite fears of disruptions to wheat production in Ukraine driven by the Russia-Ukraine war, wheat production in Ukraine in 2023 was similar to pre-war production levels. In a study of cropland abandonment in 14 regions globally, Yin et al. [18] found that detecting cropland abandonment was particularly challenging in regions characterized by small field sizes and low-intensity farms such as Nepal and Uganda. Together, these studies highlight that the impacts of armed conflict on agricultural land use is highly variable in different contexts. Quantifying changes from satellite data remains challenging, particularly in smallholder farming contexts, and requires a mixture of evidence to shed light on inaccessible regions.
[…]
@@
In this study, we used well-established cropland mapping and area estimation techniques leveraging remote sensing data, machine learning classifiers, and statistical estimators to estimate the change in cropped area in 2020 (pre-war) compared to 2021 (wartime) in Tigray. We used recommended practices for computing scientifically rigorous and transparent estimates of change in land area [20]. Our results provide evidence that complements previous studies of the impact of war on cultivation in Tigray [3, 19]. Our results show that cultivated area was largely stable between 2020 and 2021, with a relatively small amount of land attributed to cropland gain or loss. Comparing the change in cultivated area in a 5 km buffer of conflict events to the remainder of Tigray, we found a slightly higher upper limit of the confidence interval of the percentage of cropland loss within the conflict buffer (0-3%) compared to outside (0-1%). Our findings support the conclusion that Tigrayan farmers were largely able to cultivate crops at a scale comparable to pre-war years despite the widespread impact of the war on farming and other aspects of life.
@@ 2.1 Study area
[…]
@@ 2.2 Conflict data
We obtained a geographically and temporally referenced dataset of conflict events that occurred in Tigray in 2020 and 2021 from the Armed Conflict Location and Event Data Project (ACLED). ACLED compiles data from various sources including media reports, NGOs and international organizations, government reports and statements, social media and online platforms, academic literature, eyewitness accounts, and local sources. We included all event types provided in the ACLED dataset except “Peaceful Protests.” This resulted in a dataset of 831 events with multiple event types: 666 battles, 104 violence against civilians, 43 explosions/remote violence, 11 strategic developments, 5 riots, and 2 protests. Figure 1 shows a map and timeline of these events. From Figure 0(b), it can be seen that the majority of conflict events occurred prior to the start of the 2021 growing season. This plot also shows that Ethiopia’s declaration of an immediate, unilateral cease-fire occurred during the land preparation phase. However, some battles and other events were still reported after the cease-fire. Of the five admin2 zones in Tigray, the majority of the conflict events in the 2020-2021 time period were recorded in the Southern zone (247), followed by North Western (199), Central (196), Eastern (117), and Western (72).
@@ 2.3.1 Classification model
Figure 1(a) summarizes the workflow we used to create annual cropland maps. We used the long short term memory (LSTM) recurrent neural network architecture from Kerner et al. [23] to create a binary cropland map for 2020 and 2021 separately. The LSTM architecture has one base hidden layer, which is followed by two classification heads: one two-layer head is used to classify training instances within Tigray, and one single-layer head is used to classify global training instances (outside Tigray). We used a batch size of 128 instances. We used early stopping with a patience of 10 epochs and proceeded with model weights from the training checkpoint with the highest F1 score evaluated on the validation set. Training concluded after 16 epochs for the 2020 model and 13 epochs for the 2021 model (one epoch represents one complete pass through the entire training dataset). We used the Adam optimizer with default parameters to tune weights during training with learning rate set to $0.001$. We set the $\alpha$ hyperparameter that balances the global and local samples in the loss function to $10$ [23]. We used dropout during training with a dropout probability of $0.2$. We used the trained model for each year to predict the full cropland map for Tigray in 2020 and 2021 using the OpenMapFlow python library [24]. Our code used to create the cropland maps in this study can be found at https://github.com/nasaharvest/crop-mask.
@@ 2.3.2 Cropland change mapping
We created a cropland change map for 2020-2021 by combining the cropland maps created for 2020 and 2021 separately. This resulted in a map with four change classes: stable non-crop, stable crop, crop gain, crop loss.
@@ 2.3.3 Satellite data
We used the pixel-level time series input data format from the CropHarvest database [25] for the LSTM classifier. The 12-month time series consists of aggregated observations in a given pixel over each month from February 1 to January 31 for the year of interest (2020 or 2021) from Sentinel-2, Sentinel-1, and the Shuttle Radar Topographic Mission (SRTM) Digital Elevation Model (DEM). Note that while CropHarvest includes meteorological data (precipitation and temperature) from the ERA5 dataset, we omitted this dataset because we found that it introduced artifacts in the predicted map. We used all bands from Sentinel-2 except B1 (coastal aerosol) and B10 (cirrus) and added NDVI (computed from B8 and B4). This resulted in 12 channels from Sentinel-2. We aggregated the 5-day observations to monthly by taking the least-cloudy pixel per 30-day window [26]. We used top-of-atmosphere observations (Level 1A) since this is required by the algorithm. We used the VV and VH channels from the Sentinel-1 synthetic aperture radar dataset which we aggregated to monthly by taking the median of observations over each month. We used the SRTM DEM to compute elevation and slope features. In total, our input data had 16 channels for each of 12 timesteps for a given pixel (note that elevation and slope are constant over all timesteps). All datasets were pre-processed, resampled to 10 m/pixel, and exported to Google Cloud Storage using Google Earth Engine. The preprocessing code for each dataset can be found at https://github.com/nasaharvest/openmapflow/tree/main/openmapflow/eo.
2.3.4 Labeling and dataset splits
@@
Our training dataset consisted of approximately 100,000 total instances with binary labels of crop or non-crop (about 5,000 of which were in Tigray and thus used to train the local classification head). The training dataset was constructed primarily from publicly accessible data in the CropHarvest dataset (https://zenodo.org/doi/10.5281/zenodo.5021761). The validation and test datasets can be found at https://data.harvestportal.org/dataset/annual-and-change-maps-for-tigray-2020-2021. We created validation and test datasets for choosing hyperparameter settings and evaluating final performance metrics (respectively) using a random uniform sample of points in Tigray; we created independent samples for each year. Each point was labeled by the authors and other trained individuals using the Collect Earth Online platform using a combination of monthly PlanetScope (3-5 m/pixel) composites, monthly Sentinel-2 composites, Mapbox Satellite basemap, and other high-resolution images available from Google Earth. Annotators answered the question “Does this point contain active cropland?” for each point based on whether the point location visually resembled an agricultural field and showed signs of sowing, growing, and harvesting during the approximate crop calendar of Tigray. As in Skakun et al. [7], we defined cropland as “a piece of land that is sown/planted and harvestable at least once within the 12 months after the sowing/planting date.” Each point was labeled by at least two different annotators and we retained only the points for which both (or a majority of) labelers agreed. This resulted in a validation set of 440 labels (140 crop and 300 non-crop) for 2020 and 289 labels (93 crop and 196 non-crop) for 2021, and a test set of 425 labels (127 crop and 298 non-crop) for 2020 and 297 labels (109 crop and 188 non-crop) for 2021.

### v2

@@ Abstract
The Tigray War was an armed conflict that took place primarily in the Tigray region of northern Ethiopia from November 3, 2020 to November 2, 2022. Given the importance of agriculture in Tigray to livelihoods and food security, determining the impact of the war on cultivated area is critical. However, quantifying this impact was difficult due to restricted movement within and into the region and conflict-driven insecurity and blockages. Using satellite imagery and statistical area estimation techniques, we assessed changes in crop cultivation area in Tigray before and during the war. Our findings show that cultivated area was largely stable between 2020-2021 despite the widespread impacts of the war. We estimated $1,132,000\pm 133,000$ hectares of cultivation in pre-war 2020 compared to $1,217,000\pm 132,000$ hectares in wartime 2021. Comparing changes inside and outside of a 5 km buffer around conflict events, we found a slightly higher upper confidence limit of cropland loss within the buffer (0-3%) compared to outside the buffer (0-1%). Our results support other reports that despite widespread war-related disruptions, Tigrayan farmers were largely able to sustain cultivation. Our study demonstrates the capability of remote sensing combined with machine learning and statistical techniques to provide timely, transparent area estimates for monitoring food security in regions inaccessible due to conflict.
@@ 1 Introduction
Satellite remote sensing data can provide qualitative and quantitative insights about the impact of conflict on agricultural land use and production. This remote, real-time data source is especially critical for assessing impacts in regions where traditional ground-based surveys cannot be conducted due to insecurity.
Most studies using satellite data to assess war impacts on cultivation in other regions have found significant reductions in cultivated land due to war, ranging from approximately 15-30% across studies. Skakun et al. [6] found a 22% net loss of cropland in regions not occupied by the Ukraine government following the military conflict that began in 2014 in southeastern Ukraine, and that cropland losses were higher (46%) in areas in the buffer zone surrounding the conflict border. In a study of the more recent war in Ukraine that began in 2022, Ma et al. [7] detected concentrations of fallowed croplands in western Kherson and central Luhansk associated with conflict activities. Witmer [8] found that the average cultivated area decreased by 25% in the period after the Bosnian War between 1992-1995. In a study of the armed conflict in Mali, Boudinaud and Orenstein [9] found that 25% of areas in Mopti exhibited crop area losses in 2019 compared to pre-conflict years. Olsen et al. [10] concluded that cropland cultivation decreased by 16% from 2016 to 2018 as a result of armed conflict in South Sudan. In the first (1994-1996) and second (1999-2009) Chechen wars, Yin et al. [11] found higher rates of cropland abandonment in areas closer to conflict events (45% abandonment within 1 km compared to 6% greater than 8 km from conflict events) as well as higher rates of re-cultivation in areas farther away from conflict events. Müller and Kummerle [12] and Peterson and Aunap [13] studied the impact of government and economic transitions in Romania after 1989 and Estonia between 1990-1993, respectively. While these transitions were not armed conflicts, they were tumultuous periods with disruptions to the agricultural sector that were associated with 32% of croplands being abandoned in Estonia and 21% abandonment in Romania.
While most studies concluded net losses or negative impacts to agricultural production due to conflict, Li et al. [14] found that there was a net increase in rural areas of Syria during the civil war. Eklund et al. [15] reported a mixture of cropland expansion, abandonment, and other changes in areas of Syria and Iraq occupied by the Islamic State. NASA Harvest [16] found that despite fears of disruptions to wheat production in Ukraine driven by the Russia-Ukraine war, wheat production in Ukraine in 2023 was similar to pre-war production levels. In a study of cropland abandonment in 14 regions globally, Yin et al. [17] found that detecting cropland abandonment was particularly challenging in regions characterized by small field sizes and low-intensity farms such as Nepal and Uganda.
Together, these studies highlight that the impacts of armed conflict on agricultural land use is highly variable in different contexts. Quantifying changes from satellite data remains challenging, particularly in smallholder farming contexts, and requires a mixture of evidence to shed light on inaccessible regions.
@@
Agriculture is a critical industry in Tigray, providing employment for 80% of people and constituting 65% of the land area [18]. Disruptions to cultivation could have severe consequences for livelihoods and food security. Thus, quantifying the impact of the war on agriculture is crucial for targeting humanitarian relief to ensure the population’s basic needs are met. Many farmers reported delays in ploughing due to insecurity [3] and 353,000 people in Tigray were reportedly experiencing “catastrophic” levels of “acute food insecurity” in June 2021 [18]. Food insecurity in early 2021 was exacerbated by the loss of food stocks destroyed by armed forces during the start of the war in late 2020 and the blockage of food and humanitarian aid to the region [18]. Humanitarian aid was blocked or severely limited from the start of the conflict until April 2021, when aid organizations could deliver food aid for the first time since late 2020. In surveys conducted between April and May 2021, 50-70% of households in NorthWestern and Central Tigray reported that food assistance was their primary source of food [18]. The June 2021 cease-fire gave hope that farmers would be able to cultivate their fields in the 2021 main season, but there were grave concerns and uncertainty about food insecurity and the fate of food production in 2021.
Quantifying the impact of the Tigray civil war on agricultural land use is challenging due to lack of ground-truth data, small field sizes, and heterogeneous landforms. Studies of the impact of the Tigray war on cultivated area found that there was significant and widespread disruption of farming activities particularly in late 2020 and early 2021 including destruction of farm stocks and supplies, delays in ploughing and planting due to threats from soldiers, and market disruptions [19, 3, 18]. Nyssen et al. [3] reported that despite these disruptions, the effect on cultivated area was relatively moderate because most farmers were able to employ wartime adaptation strategies including shifting the timing and location of farming activities, changing crop types, and leveraging communal aid. Through remote sensing analysis, Weldegebriel et al. [19] concluded that there was a modest net loss of 714 km2 of “well cultivated land” in highland cropland areas of Tigray (which excludes the Western and parts of the North Western zones) in wartime 2021 compared to pre-war 2019/2020. These studies suggest that the impact of the war on crop cultivation in Tigray was likely moderate. However, there has not yet been a statistically rigorous estimate of the change in cultivated area over the entire region of Tigray before and after the war began.
In this study, we used well-established cropland mapping and area estimation techniques leveraging remote sensing data, machine learning classifiers, and statistical estimators to estimate the change in cropped area in 2020 (pre-war) compared to 2021 (wartime) in Tigray. We used recommended practices for computing scientifically rigorous and transparent estimates of change in land area [20]. While prior work discussed the potential for multifaceted impacts of the Tigray war on food security and the agricultural sector—such as population displacement, changes in crop types or cultivation practices, movement and access limitations, market dysfunction and loss of income, or destruction of food stocks and farming equipment [3, 18]—most of these factors are difficult or impossible to determine from satellite observations, especially in the absence of extensive ground-truth data. Our study focused on impacts of the Tigray war on crop cultivation that resulted in a binary change in cultivation status (either cultivated or not cultivated), which could be observed from satellite observations.
@@
Our results provide evidence that complements previous studies of the impact of war on cultivation in Tigray [3, 19]. Our results show that cultivated area was largely stable between 2020 and 2021, with a relatively small amount of land attributed to cropland gain or loss. Comparing the change in cultivated area in a 5 km buffer of conflict events to the remainder of Tigray, we found a slightly higher upper limit of the confidence interval of the percentage of cropland loss within the conflict buffer (0-3%) compared to outside (0-1%). Our findings support the conclusion that Tigrayan farmers were largely able to cultivate crops at a scale comparable to pre-war years despite the widespread impact of the war on farming and other aspects of life.
@@ 2.1 Study area
The Tigray region spans approximately 50,000 km2 in northern Ethiopia with an estimated population of approximately 5 million as of 2019 [21, 22]. Tigray shares a border with Eritrea to the north and Sudan to the west. Tigray contains five zones: Eastern, North Western, Western, Central, and Southern. The terrain in Tigray is very mountainous and hilly, with extensive terracing used for agriculture. Elevations range from 500 meters in the northeast to almost 4,000 meters in the southwest. The climate is generally semi-arid, with average annual rainfall between 400-800 mm concentrated in the summer months (June to September). An estimated 80% of the population works in agriculture and around 65% of the land area is devoted to cultivation [21]. The Central and Western zones have more intensive agriculture due to higher rainfall and fertile soils. The majority of crop cultivation in Tigray is rainfed with two growing seasons: meher (June-September), which is the main growing season, and belg (February-May), which is the minor growing season. The majority of agricultural land in Tigray is used to grow teff, barley, wheat, maize, sorghum, finger millet, oats, and sesame [23]. For most crops, planting occurs between May to July, vegetative growth between July and September, and harvesting between September and December [21].
@@ 2.2 Conflict data
We obtained a geographically and temporally referenced dataset of conflict events that occurred in Tigray in 2020 and 2021 from the Armed Conflict Location and Event Data Project (ACLED). ACLED compiles data from various sources including media reports, NGOs and international organizations, government reports and statements, social media and online platforms, academic literature, eyewitness accounts, and local sources. We included all event types provided in the ACLED dataset except “Peaceful Protests.” This resulted in a dataset of 831 events with multiple event types: 666 battles, 104 violence against civilians, 43 explosions/remote violence, 11 strategic developments, 5 riots, and 2 protests.
Figure 1 shows a map and timeline of these events. From Figure 1(b), it can be seen that the majority of conflict events occurred prior to the start of the 2021 main growing season. This plot also shows that Ethiopia’s declaration of an immediate, unilateral cease-fire occurred during the land preparation phase. However, some battles and other events were still reported after the cease-fire. Of the five zones in Tigray, the majority of the conflict events in the 2020-2021 time period were recorded in the Southern zone (247), followed by the North Western (199), Central (196), Eastern (117), and Western (72) zones.
November 2020 marked the official start of the war and was the month with the most recorded conflict events during the study period. This also coincided with the harvest period of the main meher season of 2020. Despite this outbreak, final crop yields estimated at the regional and national level were reportedly favorable, with the exception of localized conflict-affected areas which were reported to have lost an estimated 90% of the main 2020 season harvest due to pillaging and burning by armed forces [24]. Based on the conflict timeline and previous reports (e.g., [3, 24]), we expected that the conflict might have some impact on the minor belg 2021 season but were uncertain about the likelihood of impact on the major meher 2021 season. In this study, we did not differentiate between estimated area of crops cultivated during the belg versus meher season since our goal was to estimate the change in the total cultivated between 2020 and 2021. Furthermore, estimating changes between the individual seasons would be difficult since farmers may plant the same fields in different seasons each year.
@@ 2.3.1 Classification model
Figure 2(a) summarizes the workflow we used to create annual cropland maps for 2020 and 2021. We used the long short term memory (LSTM) recurrent neural network architecture from Kerner et al. [25] to create a binary cropland map for 2020 and 2021 separately. The LSTM architecture has one base hidden layer, which is followed by two classification heads: one two-layer head is used to classify training instances within Tigray, and one single-layer head is used to classify global training instances (outside of Tigray). We used a batch size of 128 instances. We used early stopping with a patience of 10 epochs and proceeded with model weights from the training checkpoint with the highest F1 score evaluated on the validation set. Training concluded after 16 epochs for the 2020 model and 13 epochs for the 2021 model (one epoch represents one complete pass through the entire training dataset). We used the Adam optimizer with default parameters to tune weights during training with learning rate set to $0.001$. We set the $\alpha$ hyperparameter that balances the global and local samples in the loss function to $10$ [25]. We used dropout during training with a dropout probability of $0.2$. We used the trained model for each year to predict the full cropland map for Tigray in 2020 and 2021 using the OpenMapFlow python library [26]. Our code used to create the cropland maps in this study can be found at https://github.com/nasaharvest/crop-mask.
@@ 2.3.2 Cropland change mapping
We created a cropland change map for 2020-2021 by combining the cropland maps created for 2020 and 2021 separately. This resulted in a map with four change classes: stable non-crop, stable crop, crop gain, crop loss. Using the annual cropland maps to create a change map, rather than training a model to classify four change classes directly, allowed us to leverage a large training dataset available for binary cropland classification (see Section 2.3.4). Since large labeled datasets for four-class cropland change are not available, direct change classification would have required us to label a large dataset of training samples for the change classes, which would have been extremely time-consuming.
@@ 2.3.3 Satellite data
We used the pixel-level time series input data format from the CropHarvest database [27] for the LSTM classifier. The 12-month time series consists of aggregated observations in a given pixel over each month from February 1 to January 31 for the year of interest (2020 or 2021) to allow the model to observe the full growing cycle in either the belg or meher season. We included observations from Sentinel-2, Sentinel-1, and the Shuttle Radar Topographic Mission (SRTM) Digital Elevation Model (DEM). Note that while CropHarvest includes meteorological data (precipitation and temperature) from the ERA5 dataset, we omitted this dataset because we found that it introduced artifacts in the predicted map. We used all bands from Sentinel-2 except B1 (coastal aerosol) and B10 (cirrus) and added NDVI (computed from B8 and B4). This resulted in 12 channels from Sentinel-2. We aggregated the 5-day observations to monthly by taking the least-cloudy pixel per 30-day window using the method described by Schmitt et al. (2019) [28]. We used top-of-atmosphere observations (Level 1A) since this was required by the cloud-filtering algorithm. We used the VV and VH channels from the Sentinel-1 synthetic aperture radar dataset which we aggregated to monthly by taking the median of observations over each month. We used the SRTM DEM to compute elevation and slope features. In total, our input data had 16 channels for each of 12 timesteps for a given pixel (note that elevation and slope are constant over all timesteps). All datasets were pre-processed, resampled to 10 m/pixel, and exported to Google Cloud Storage using Google Earth Engine. The preprocessing code for each dataset can be found at https://github.com/nasaharvest/openmapflow/tree/main/openmapflow/eo.
2.3.4 Model training, validation, and test data
Our training dataset consisted of approximately 100,000 total instances with binary labels of crop or non-crop (about 5,000 of which were in Tigray and thus used to train the local classification head). The training dataset was constructed primarily from publicly accessible data in the CropHarvest dataset (https://zenodo.org/doi/10.5281/zenodo.5021761). Since our training data contained a mix of global and local (within Tigray) samples, we do not expect our model to be biased toward detecting cropland in a specific season.
@@ 2.3.4 Model training, validation, and test data
The validation and test datasets can be found at https://data.harvestportal.org/dataset/annual-and-change-maps-for-tigray-2020-2021. We created validation and test datasets for choosing hyperparameter settings and evaluating final performance metrics (respectively) using a random uniform sample of points in Tigray. Since points were randomly sampled within Tigray, we expect the proportion of cropland in this sample to represent the approximate proportion of cropland area in the region, including the approximate proportion of planting in the meher and belg seasons typically observed in the region. We created independent samples for each year. Each point was labeled by the authors and other trained individuals. Annotators answered the question “Does this point contain active cropland?” for each point. We describe our labeling procedure used to answer this question in Section 2.5.2. Each point was labeled by at least two different annotators and we retained only the points for which both (or a majority of) labelers agreed. This resulted in a validation set of 440 labels (140 crop and 300 non-crop) for 2020 and 289 labels (93 crop and 196 non-crop) for 2021, and a test set of 425 labels (127 crop and 298 non-crop) for 2020 and 297 labels (109 crop and 188 non-crop) for 2021.